A rough strip of negligible mass and length 10cm b
Answer #1

The moment the strip starts, a constant torque starts acting on the wheels whose magnitude can be given by

τ = Fr = (20)(1/100)= 0.2 N-m

it acts till the whole strip of length 10 cm is pulled out that means it acts till the wheels rotate through ϴ = l/r = 10/1 = 10 rad

hence the work done = 2(τϴ)= 2(0.2)(10) = 4 J (the factor of 2 comes due to two wheels)

now as initially the wheels were at rest and now they are rotating hence the change in kinetic energy = amount of work done

2(Iω2/2) = 4

2 = 4 where I = mr2/2

Now

Part a

The work done in pulling the strip = 4 J

Part b

The rotational kinetic energy of each wheel, E = Iω2/2 = 2 J

Part c

Angular momentum = (2IE)1/2 (using relation E= L2/2I)

Part d

Clearly the required angle = 10 rad

Part e

As

2 = 4

ω= (4/I)1/2

(note we need mass or moment of inertia of the wheels in order to complete the part c and e)